# Almost Surely Safe Alignment of Large Language Models at Inference-Time

### v2

@@ 4 InferenceGuard: Safety at Test-Time
To fully define our critic, we require a representation of the embedding of our augmented state $\tilde{\mathbf{s}}_{t}=[\mathbf{s}_{t},\text{z}_{t}]$ within the latent space. As noted above, we can acquire $(\mathbf{h}_{t},\mathbf{o}_{t})$ from the transformer architecture. We call this mapping $\phi$, whereby $(\mathbf{h}_{t},\mathbf{o}_{t})=\phi(\{\mathbf{x},\mathbf{y}_{<t}\})$. Furthermore, we use an identity mapping to embed $\text{z}_{t}$, which enables us to input the actual tracking of the constraints directly to the critic without any loss of information.
@@ 4.1 Theoretical Insights
In this section, we show that optimizing in the latent space preserves safety constraints in the original token space, and we prove that our approach guarantees almost sure safety.
We should answer two essential questions: i) Can we still compute an optimal policy in the latent space? and ii) If we enforce safety constraints in the latent space, do they still hold in the original token space? While the prior work in (Sootla et al., 2022) established theoretical results for safety-augmented MDPs in standard (non-LLM) RL settings, their work does not address how guarantees in the latent space translate to the original token space. To handle those problems, we extend the theorems from Hernández-Lerma & Muñoz de Ozak (1992); Sootla et al. (2022) to ensure the following three properties:
@@
  • Prop II) Policies and value functions in the latent space are valid in the original token space, implying that optimizing in the latent space preserves the constraints in original token space (Theorem 4.1 (b,c))
@@
where $\tilde{\mathbf{s}}_{t}$ is the augmented state in the original token space.
According to Definition 4.1, we ensure that the cost incurred by the augmented state $\tilde{\mathbf{s}}_{t}=[\{\mathbf{x},\mathbf{y}_{<t}\},\text{z}_{t}]$ w.r.t. $\tilde{\mathcal{C}}^{n}_{\text{task}}$ is equal to the latent cost incurred by the latent state $[\phi(\{\mathbf{x},\mathbf{y}_{<t}\}),\text{z}_{t}]$ w.r.t $\bar{\mathcal{C}}_{\text{task}}^{n}$. Additionally, we ensure that the transition dynamics of the augmented state in the original token space and the corresponding latent state in the latent space are equivalent.
This equivalence enables us to derive an optimal policy for the latent MDP and apply it to minimize the cost objective in the original augmented MDP (see Equation 7). We proceed to analyze the existence of such an optimal policy in the latent space through the following standard assumptions (Sootla et al., 2022) on $\bar{\mathcal{C}}^{n}_{\text{task}}$, and $\bar{\mathcal{P}}$: A1. The function $\bar{\mathcal{C}}_{\text{task}}^{n}(\mathbf{h},\mathbf{o},\text{z},\text{y})$ is bounded, measurable, nonnegative, lower semi-continuous w.r.t. $(\mathbf{h},\mathbf{o},\text{z})$ for a given y, and A2. The transition law $\bar{{\mathcal{P}}}$ is weakly continuous for any y.
Next, we define $\bar{\pi}$ as a policy in the latent space that maps $(\mathbf{h},\mathbf{o},\text{z})\rightarrow\text{y}$, and its value function for an initial state $(\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})$ as follows: $\bar{V}^{n}(\bar{\pi},\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})=\mathbb{E}_{\bar{\pi}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}_{\text{task}}^{n}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z},\text{y})\Big{]}$. Then, one can define the optimal value function:
@@
Since we cannot optimize directly in the original constrained MDP, we first show that solving for an optimal policy in the latent MDP preserves key properties of the original problem. The following theorem formalizes this by proving the existence of the optimal policy and its mapping to the original MDP. The proof is in Appendix B.3.
@@
thmthmsauteequivalence(Optimality in the Latent Space) Given the latent MDP in Definition 4.1 and with A1-A2, we can show:
@@
    (Prop I) For any finite $n$, the Bellman equation holds, i.e., there exists a function $\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})$ such that:
    $\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})=\min_{\text{y}\in{\mathcal{V}}}\Big{(}\bar{\mathcal{C}}^{n}_{\text{task}}(\mathbf{h},\mathbf{o},\text{z},\text{y})$
    $\hskip 130.0002pt+\gamma\bar{V}^{\star,n}(\mathbf{h}^{\prime},\mathbf{o}^{\prime},\text{z}^{\prime})\Big{)},$
    $\text{such that}\ \ (\mathbf{h}^{\prime},\mathbf{o}^{\prime},\text{z}^{\prime})\sim\bar{{\mathcal{P}}}(\cdot|\mathbf{h},\mathbf{o},\text{z},\text{y})$
@@
The above theorem ensures that finding and securing the existence of the optimal policy in the latent space is sufficient to solve Equation 7 optimally [^5] As noted in Section 3.2, we analyze the transition from a large $n$ to $\infty$, and confirm that the results hold even for $\tilde{\mathcal{C}}_{\text{task}}^{\infty}$.. Informally, the latent space acts as a faithful representation, preserving constraints and making optimization computationally efficient. This implies that the optimal policies and value functions in the latent space remain valid in the original space.
Almost Sure Guarantee. Now, we derive Prop III that ensures the safety cost constraints are almost surely satisfied. This is more challenging than Equation 3, where only the expected safety cost is constrained:
$\min_{\pi}$ $\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t})\Big{]}$ (9)
@@ 4.2 Algorithm and Practical Implementation
Training a Latent-Space Critic. We make the usual assumption that trajectories terminate at a maximum length $T$. In this case, the value function simplifies to become: $\bar{V}^{n}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t})=\mathbb{E}_{\bar{\pi}}[\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})]$ if there is safety budget left, i.e., if $\text{z}_{T}>0$, or $n$ if $\text{z}_{T}\leq 0$, where $\bar{c}_{\text{task}}({\mathbf{h}}_{t},{\mathbf{o}}_{t})=c_{\text{task}}([\mathbf{x},\mathbf{y}_{\leq t}])$ in the latent MDP.
@@
Hence, it is sufficient to predict: the sign of $\text{z}_{T}$ and the value of $\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})$ to assess the quality of a state. We estimate those through Monte Carlo (MC) sampling. Specifically, we generate multiple trajectories from the initial state ($\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})$ using the reference policy, and compute the mean terminal cost, the sign of $z_{T}$ to serve as targets for the critic training. The usual alternative to MC sampling is Temporal Difference (TD) learning, where the critic is updated based on the difference between the current estimate and a bootstrapped estimate from the next state. However, MC sampling offers two advantages: i) it simplifies training by using separate supervised signals for quality and safety, unlike TD, which combines both, and ii) it allows dynamic adjustment of $n$ without retraining.
@@
We train a critic network with two heads by sampling responses from the base model and scoring them using the cost function. We define $\mathcal{J}_{1}$ as the binary cross-entropy for predicting the sign of $\text{z}_{T}$ and $\mathcal{J}_{2}$ as the mean squared error for predicting $\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})$. Our critic training minimizes: $\mathcal{J}(\mathbf{\theta})=\mathbb{E}_{\bar{\pi}}\Big{[}\sum_{t=1}^{T}\mathcal{J}_{1}\left(f^{1}_{\mathbf{\theta}}({\mathbf{h}}_{t},{\mathbf{o}}_{t},\text{z}_{t}),\text{z}_{T}>0\right)+\mathcal{J}_{2}\left(f^{2}_{\mathbf{\theta}}({\mathbf{h}}_{t},{\mathbf{o}}_{t},\text{z}_{t}),\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})\right)\Big{]}$, where $f^{1}_{\mathbf{\theta}}$ and $f^{2}_{\mathbf{\theta}}$ are the heads of our parameterized critic.
@@
Search method. We build on the beam search strategy from (Mudgal et al., 2023; Li, 2025) wherein we sequentially sample $N$ beams of $d$ tokens into a set $B$ from the pre-trained model and choose $K$ beams with the highest scores as possible continuations of the prompt (see Algorithm 1 in Appendix C). This step ensures that we focus on the most promising continuations. The goal of the scoring function is to balance the immediate task cost and the predicted future task cost while ensuring safety. This is repeated until we complete trajectories. Given a token trajectory $\text{y}_{t:t+d}$, we present a scoring function $\text{E}_{\text{critic}}$ where we assume that we cannot evaluate intermediate answers with the cost functions. However, when immediate safety costs are available, a simpler scoring function, see Appendix C. We define $\text{E}_{\text{critic}}$ as:
@@
Figure 2: Trade-offs between safety, reward, and inference time for BoN, ARGS, RECONTROL, Beam Search, and InferenceGuard on the PKU-SafeRLHF test dataset, evaluated for Alpaca-7B (top) and Beaver-v3-7B (bottom). Reward is the average score evaluated by the reward model, safety rate is the percentage of tasks completed within budget $d$, and inference time is the average per-task duration. The left column (reward vs. safety) and right column (inference time vs. safety) categorize methods by performance. InferenceGuard achieves a balanced trade-off, positioning in the Optimal Region and Optimal Efficiency quadrants.
Sampling Diversity. Finally, if the right selection strategy can guarantee that we will converge on a safe solution, it does not consider how many samples would be necessary. To increase the search speed, we introduce a diversity term in the sampling distribution when no safe samples were found based on the token frequency of failed beams. We denote $F$ as the frequency matrix counting the tokens we previously sampled from $t$ to $t+d$. Instead of sampling from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t})$, we sample from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t}-\text{n}_{2}({F_{t}>0}))$ where $\text{n}_{2}({F_{t}>0})$ is a vector where each component $j$ is $\text{n}_{2}$ if $F_{t,j}>0$ and 0 otherwise. The addition of $\text{n}_{2}({F_{t}>0})$ disables the possibility of sampling the same token at the same position observed in unsuccessful beams, thus increasing diversity.
It is worth noting that as we sample from the reference LLM and rank responses directly or via the critic, block sampling ensures a small Kullback-Leibler (KL) divergence from the original LLM without explicitly adding a KL regularizer into our objective, preserving coherence and natural flow; see (Mudgal et al., 2023).
@@ 5 Experiments
Figure 3: Reward and cost distributions on PKU-SafeRLHF test tasks using Alpaca-7B (top) and Beaver-v3 (bottom) as base models. The left y-axis shows reward distribution. The right y-axis shows the maximum cumulative cost. InferenceGuard outperforms others, achieving higher rewards with lower costs across both models.
@@
Baselines. We evaluate the helpfulness (task cost) and harmlessness (safety cost) of our method on both non-safety-aligned and safety-aligned base models, Alpaca-7B Taori et al. (2023) and Beaver-7B (Ji et al., 2024b). We compare InferenceGuard to the following state-of-the-art test-time alignment methods: Best-of-N (BoN), beam search, ARGS (Khanov et al., 2024) and RECONTROL (Kong et al., 2024). These test-time alignment methods were originally designed to maximize rewards without considering safety. To ensure a fair and meaningful comparison, we extend them to incorporate safe alignment for LLMs. We incorporate a Lagrangian-based approach for all baselines that penalizes safety violations, ensuring a balanced trade-off between task performance and safety. This helps us evaluate our performance against other algorithms and highlights the importance of safety augmentation instead of a Lagrangian approach for effectively balancing rewards and constraint satisfaction.
@@
For beam search and Best-of-N (BON), we adopt a Lagrangian approach to select solutions with $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ where $\lambda$ is the Lagrangian multiplier. Similarly, we extend ARGS so that token selection follows: $-\omega\pi(t|\cdot)+c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$, with $\omega$ adjusting the influence of the reference policy. We also considered state augmentation for ARGS and RE-control but found it ineffective. Since these methods decode token-by-token, they cannot recover once $\text{z}_{t}$ flips the sign, and before that, $\text{z}_{t}$ has no influence. Thus, we excluded it from our evaluation. To further strengthen our comparison, we introduce safety-augmented versions of BoN and beam search as additional baselines.
Datasets. We evaluate all the above methods on the PKU-SafeRLHF dataset (Ji et al., 2024b), a widely recognized benchmark for safety assessment in the LLM literature. This dataset contains 37,400 training samples and 3,400 testing samples. Of course, our training samples are only used to train the critic value network. Here, we construct a dataset by generating responses using the base models with five sampled trajectories for each PKU-SafeRLHF prompt from the training set.
@@
Evaluation Metrics. We assess the performance using several metrics: the Average Reward is computed using the reward model of (Khanov et al., 2024) as $-c_{\text{task}}$ on the complete response to reflect helpfulness, where a higher reward indicates better helpfulness; the Average Cost is evaluated with the token-wise cost model from (Dai et al., 2023) as $\mathcal{C}_{\text{safety}}$, indicating harmfulness, with higher cost values reflecting more resource-intensive or harmful outputs; the Safety Rate is defined as the proportion of responses where the cumulative cost does not exceed the safety budget $z_{t=0}=10$, and is formally given by Safety $\text{Safety Rate}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\mathcal{C}^{i}_{\text{test}}\leq z_{t=0})$, where $N$ is the total number of responses; and Time refers to the inference time taken to generate a response in seconds.
@@
Results. We present our main results in Figure 2 and additional ones in Table 2 in Appendix D. InferenceGuard achieves the highest safety rates with both models (91.04% and 100% for Alpaca and Beaver respectively). With Beaver, our method dominates the Patero front, achieving the highest rewards without any unsafe responses. Although Lagrangian methods can have a reasonable average cost, they all fail to satisfy the safety constraints. In the meantime, they are too safe on already safe answers, hindering their rewards. The RE-control intervention method underperforms for both models in our setting. ARGS can provide safe answers but with very poor rewards because most answers are very short to avoid breaking the safety constraint. Best-of-N and beam search with augmented safety came after us. However, Best-of-N cannot leverage intermediate signal, and beam search is blind to its previous mistakes, yielding more unsafe answers.
@@
Figure 3 provides a better view of the reward and safety distributions. The figure shows that InferenceGuard consistently achieves higher rewards while maintaining low cumulative costs, outperforming other methods across both models. Its maximum cumulative cost stays just under the safety budget while maximizing the reward as suggested by our theoretical contributions. Finally, we observe that the trained critic helps to better guide the search on intermediate trajectories in both settings. We also compare generated answers in Appendix D qualitatively.
6 Related Work
Safe RL: Safe RL employs the cMDP framework, formulated in (Altman, 1999). Without prior knowledge, (Turchetta et al., 2016; Koller et al., 2018; Dalal et al., 2018) focus on safe exploration, while with prior knowledge, (Chow et al., 2018; 2019; Berkenkamp et al., 2017) learn safe policies using control techniques. (Achiam et al., 2017; Ray et al., 2019; Stooke et al., 2020) enforce safety constraints via Lagrangian or constrained optimization which can often lead to suboptimal safety-reward trade-offs. Instead, we extend safety state augmentation Sootla et al. (2022) to LLMs and latent MDPs to ensure almost sure inference time safety.
LLM alignment and safety: Pre-trained LLMs are often aligned to specific tasks using RL from Human Feedback (RLHF), where LLMs are fine-tuned with a learned reward model (Stiennon et al., 2020; Ziegler et al., 2019; Ouyang et al., 2022) or directly optimized from human preferences (Rafailov et al., 2023; Azar et al., 2023; Zhao et al., 2023; Tang et al., 2024; Song et al., 2024; Ethayarajh et al., 2024). (Bai et al., 2022; Ganguli et al., 2022) first applied fine-tuning in the context of safety, and (Dai et al., 2023) proposed safe fine-tuning via Lagrangian optimization. Other safety-focused methods such as (Gundavarapu et al., 2024; Gou et al., 2024; Hammoud et al., 2024; Hua et al., 2024; Zhang et al., 2024b; Guo et al., 2024; Xu et al., 2024; Wei et al., 2024; Li et al., 2025a) are either orthogonal, handle a different problem to ours, or can not ensure almost sure safety during inference.
Inference time alignment: Inference-time alignment of LLMs is often performed through guided decoding, which steers token generation based on rewards (Khanov et al., 2024; Shi et al., 2024; Huang et al., 2024) or a trained value function Han et al. (2024); Mudgal et al. (2023); Kong et al. (2024). Other safety-focused inference-time methods include (Zhong et al., 2024b; Banerjee et al., 2024; Niu et al., 2024; Wang et al., 2024a; Zeng et al., 2024; Zhao et al., 2024). Compared to those methods, we are the first to theoretically guarantee almost sure safe alignment with strong empirical results. Operating in the latent space enables us to train smaller, inference-efficient critics while optimally balancing rewards and safety constraints without introducing extra parameters, e.g., Lagrangian multipliers.
We detail other related works extensively in Appendix A.

### v3

@@ 5 InferenceGuard: Safety at Test-Time
To define our critic, we require a representation of our augmented state $\tilde{\mathbf{s}}_{t}=[\mathbf{s}_{t},\text{z}_{t}]$ within the latent space. As noted above, we can acquire $(\mathbf{h}_{t},\mathbf{o}_{t})$ from the transformer architecture. We call this mapping $\phi$, whereby $(\mathbf{h}_{t},\mathbf{o}_{t})=\phi(\{\mathbf{x},\mathbf{y}_{<t}\})$. To embed $\text{z}_{t}$, we use an identity mapping which enables us to input the actual tracking of the constraints directly to the critic without any loss of information.
@@ 5.1 Theoretical Insights
In this section, we show that optimizing in the latent space preserves safety constraints in the original token space and prove that our approach guarantees almost sure safety w.r.t. given safety cost model.
We consider two essential questions: i) Can we compute an optimal policy in the latent space?, ii) If we enforce safety constraints in the latent space, do they still hold in the original token space? While [16] established theoretical results for safety-augmented MDPs in standard (non-LLM) RL settings, their work does not address how guarantees in the latent space translate to the original token space. To handle those problems, we extend the theorems from [15, 16] to ensure the following properties:
@@
  • Prop II) The latent space policies and value functions are valid in the original token space. Hence, optimizing in the latent space preserves the constraints in original token space (Theorem 5.1 (b,c))
@@
where $\tilde{\mathbf{s}}_{t}$ is the augmented state in the original token space. Definition 5.1 ensures that the cost incurred by the augmented state $\tilde{\mathbf{s}}_{t}=[\{\mathbf{x},\mathbf{y}_{<t}\},\text{z}_{t}]$ w.r.t. $\tilde{\mathcal{C}}^{n}_{\text{task}}$ is equal to the latent cost incurred by the latent state $[\phi(\{\mathbf{x},\mathbf{y}_{<t}\}),\text{z}_{t}]$ w.r.t $\bar{\mathcal{C}}_{\text{task}}^{n}$. Moreover, it ensures that the transition dynamics of the augmented state in the original token space and the corresponding latent state in the latent space are equivalent.
This equivalence enables us to derive an optimal policy for the latent MDP and apply it to minimize the cost objective in the original augmented MDP (see Equation 7). We proceed to analyze the existence of such an optimal policy in the latent space through the following standard assumptions [16] on $\bar{\mathcal{C}}^{n}_{\text{task}}$, and $\bar{\mathcal{P}}$: A1. The function $\bar{\mathcal{C}}_{\text{task}}^{n}(\mathbf{h},\mathbf{o},\text{z},\text{y})$ is bounded, measurable, nonnegative, lower semi-continuous w.r.t. $(\mathbf{h},\mathbf{o},\text{z})$ for a given y, and A2. The transition law $\bar{{\mathcal{P}}}$ is weakly continuous for any y.
Next, we define $\bar{\pi}$ as a policy in the latent space that maps $(\mathbf{h},\mathbf{o},\text{z})\rightarrow\text{y}$, and its value function for an initial state $(\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})$ as follows: $\bar{V}^{n}(\bar{\pi},\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})=\mathbb{E}_{\bar{\pi}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}_{\text{task}}^{n}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t},\text{y})\Big{]}$. Then, one can define the optimal value function:
@@
Since we cannot optimize directly in the original constrained MDP, we first show that solving for an optimal policy in the latent MDP preserves key properties of the original problem. The following theorem formalizes this by proving the existence of the optimal policy and its mapping to the original MDP.
@@
thmthmsauteequivalence(Optimality in the Latent Space) Given A1-A2, the latent MDP in Definition 5.1 satisfies:
@@
    (Prop I) For any finite $n$, the Bellman equation holds, i.e., there exists $\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})$ such that:
    $\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})=\min_{\text{y}\in{\mathcal{V}}}\Big{(}\bar{\mathcal{C}}^{n}_{\text{task}}(\mathbf{h},\mathbf{o},\text{z},\text{y})+\gamma\bar{V}^{\star,n}(\mathbf{h}^{\prime},\mathbf{o}^{\prime},\text{z}^{\prime})\Big{)},(\mathbf{h}^{\prime},\mathbf{o}^{\prime},\text{z}^{\prime})\sim\bar{{\mathcal{P}}}(\cdot|\mathbf{h},\mathbf{o},\text{z},\text{y})$
@@
The above theorem ensures that finding and securing the existence of the optimal policy in the latent space is sufficient to solve Equation 7 optimally. Informally, the latent space acts as a faithful representation, preserving constraints and making optimization computationally efficient. This implies that the optimal policies and value functions in the latent space remain valid in the original space. We relegate the proof to Appendix B.3 but outline the main steps for Prop II below:
We first show the infimum operator recovers lower semi continuity, which implies for finite $n$ ,$\bar{V}^{\star,n}$ is lower semicontinous due to assumptions A1, A2. Next, we formulate $\bar{V}^{\star,n}\to\bar{V}^{\star,\infty}$, as the convergence of limit infimum of a sequence of increasing lower semi continuous functions, $v_{n}\to v_{0}$, to infimum of $v_{0}$. We construct a decreasing sequence of compact action sets consisting of the infimum actions $a_{n}$ of $v_{n}$ and converging to the infimum action set of $v_{0}$ consisting of $a_{0}$. We use properties of the discrete action space to show that this sequence of actions has a convergent subsequence $a_{n_{i}}\to a_{0}$ and use that to show that limit infimum of $v_{n}$ converges to infimum of $v_{0}$ (see part-c of Lemma 1 ).
Now, we derive Prop III that ensures the safety cost constraints are almost surely satisfied. This is more challenging than Equation 3, where only the expected safety cost is constrained:
@@
Theorem 5.1 justifies our state-augmented MDP reformulation, as it guarantees that a solution to the augmented MDP (Equation 7) almost surely satisfies the constraints of Equation 3 distinguishing its effectiveness from alternative reformulations that may not offer such a guarantee.
@@ 5.2 Algorithm and Practical Implementation
Training a Latent-Space Critic. We make the usual assumption that trajectories terminate at a maximum length $T$. In this case, the value function simplifies to become: $\bar{V}^{n}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t})=\mathbb{E}_{\bar{\pi}}[\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})]$ if there is safety budget left, i.e., if $\text{z}_{T}>0$, or $n$ if $\text{z}_{T}\leq 0$, where $\bar{c}_{\text{task}}({\mathbf{h}}_{t},{\mathbf{o}}_{t})=c_{\text{task}}([\mathbf{x},\mathbf{y}_{\leq t}])$ in the latent MDP. Hence, it is sufficient to predict: the sign of $\text{z}_{T}$ and the value of $\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})$ to assess the quality of a state. We estimate those through Monte Carlo (MC) sampling. Specifically, we generate multiple trajectories from the initial state ($\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})$ using the reference policy, and compute the mean terminal cost, the sign of $z_{T}$ to serve as targets for the critic training. The usual alternative to MC sampling is Temporal Difference (TD) learning, where the critic is updated based on the difference between the current estimate and a bootstrapped estimate from the next state. However, MC sampling offers two advantages: i) it simplifies training by using separate supervised signals for quality and safety, unlike TD, which combines both, and ii) it allows dynamic adjustment of $n$ without retraining.
@@
We train a critic network with two heads ( $f^{1}_{\mathbf{\theta}}$ and $f^{2}_{\mathbf{\theta}}$) by sampling responses from the base model and scoring them using the cost function. We define $\mathcal{J}_{1}$ as the binary cross-entropy for predicting the sign of $\text{z}_{T}$ and $\mathcal{J}_{2}$ as the mean squared error for predicting $\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})$. Our critic training minimizes: $\mathcal{J}(\mathbf{\theta})=\mathbb{E}_{\bar{\pi}}\Big{[}\sum_{t=1}^{T}\mathcal{J}_{1}\left(f^{1}_{\mathbf{\theta}}({\mathbf{h}}_{t},{\mathbf{o}}_{t},\text{z}_{t}),\text{z}_{T}>0\right)+\mathcal{J}_{2}\left(f^{2}_{\mathbf{\theta}}({\mathbf{h}}_{t},{\mathbf{o}}_{t},\text{z}_{t}),\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})\right)\Big{]}$.
@@
Search method. We build on the beam search strategy from [14, 66] wherein we sequentially sample $N$ beams of $d$ tokens from the pre-trained model and choose $K$ beams with the highest scores as possible continuations of the prompt (see Algorithm 1 in Appendix C). This ensures that we focus on the most promising continuations. The goal of the scoring function is to balance the immediate task cost and the predicted future task cost while ensuring safety. This is repeated until we complete trajectories. Given a token trajectory $\text{y}_{t:t+d}$, we present a scoring function $\text{E}_{\text{critic}}$ that assumes we cannot evaluate intermediate answers with the cost functions. However, when immediate safety costs are available, a simpler scoring function can be used, see Appendix C. We define $\text{E}_{\text{critic}}$ as:
@@
Figure 2: Trade-offs between safety, reward, and inference time evaluated on Alpaca-7B and Beaver-v3-7B using the PKU-SafeRLHF dataset (top), and Vicuna-7B and LLaMA3-8B using the HEx-PHI and HH-RLHF datasets respectively (bottom). Each subplot visualizes the average reward score versus safety rate, and the inference time denoted by the color of each symbol. InferenceGuard achieves a well-balanced trade-off across all three objectives by positioning in the Optimal Region.
Sampling Diversity. Finally, if the right selection strategy can guarantee that we will converge on a safe solution, it does not consider how many samples would be necessary. To increase the search speed, we introduce a diversity term in the sampling distribution when no safe samples were found based on the token frequency of failed beams. We denote $F$ as the frequency matrix counting the tokens we previously sampled from $t$ to $t+d$. For each step $i\in[1,d]$, instead of sampling from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t+i})$, we resample from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t+i}-\text{n}_{2}({F_{i}>0}))$ where $\text{n}_{2}({F_{i}>0})$ is a vector where each component $j$ is $\text{n}_{2}$ if $F_{i,j}>0$ and 0 otherwise. The addition of $\text{n}_{2}({F_{i}>0})$ disables the possibility of sampling the same token at the same position observed in unsuccessful beams, thus increasing diversity.
It is worth noting that as we sample from the reference LLM and rank responses directly or via the critic, block sampling ensures a small Kullback-Leibler (KL) divergence from the original LLM without explicitly adding a KL regularizer into our objective, preserving coherence and flow; see [14].
@@ 6 Experiments
Baselines. We evaluate the helpfulness (task cost) and harmlessness (safety cost) of our method on four models with varying safety alignment levels: the Alpaca-7B model [67], Vicuna-7B [68], Llama3.1-8B-Instruct [69] and the safety-aligned Beaver-v3-7B model [70]. We compare InferenceGuard to the following state-of-the-art test-time alignment methods: Best-of-N (BoN), beam search, and more recent advances ARGS [43] and RE-Control [42] that combines token probabilities with reward scores predicted by pre-trained reward model or value network during the decoding process. These test-time alignment methods were originally designed to maximize rewards without considering safety. To ensure a fair and meaningful comparison, we extend them to also align for safety through the Lagrangian-based approach and the safety augmentation approach. This helps us evaluate our performance against other algorithms and highlights the importance of safety augmentation over Lagrangian approach for effectively balancing rewards and constraint satisfaction.
@@
For beam search and Best-of-N (BON), we select solutions with $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ where $\lambda$ is the Lagrangian multiplier. Similarly, we extend ARGS so that token selection follows: $-\omega\pi(t|\cdot)+c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$, with $\omega$ adjusting the influence of the reference policy. We also considered state augmentation for ARGS and RE-Control but found it ineffective. Since these methods decode token-by-token, they cannot recover once $\text{z}_{t}$ flips the sign, and before that, $\text{z}_{t}$ has no influence. Thus, we excluded it from our evaluation. To further strengthen our comparison, we introduce safety-augmented versions of BoN and beam search as additional baselines.
Datasets. We evaluate across three widely recognized safety assessment benchmarks with varying sensitivity: 1) PKU-SafeRLHF includes 37,400 training samples and 3,400 testing samples of safety-critical instructions for reward and cost alignment; 2) HEx-PHI consists of 330 harmful instructions across 11 safety-relevant categories; and 3) HH-RLHF contains 112,000 training samples and 12,500 testing samples of high-safety prompts with strong human preferences. To train the critic network, we construct a dataset by generating five responses per training set prompt from the base model. For HEx-PHI evaluation, we use a value network trained on HH-RLHF, due to the limited size of HEx-PHI data.
@@
Evaluation Metrics. We assess the performance using several metrics: the Average Reward is computed using the reward models from [43] for Llama2-based models and [71] for the Llama3-based model as $-c_{\text{task}}$ on the complete response to reflect helpfulness, where a higher reward indicates better helpfulness; the Average Cost is evaluated with the cost model from [28] and [71], for Llama2 and Llama3 models respectively, as $\mathcal{C}_{\text{safety}}$, indicating harmfulness, with higher cost values reflecting more harmful outputs; the Safety Rate is the proportion of responses where the cumulative cost does not exceed the safety budget $z_{t=0}=10$, and is given by Safety $\text{Safety Rate}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\mathcal{C}^{i}_{\text{test}}\leq z_{t=0})$, where $N$ is the number of prompts; and Inference Time refers to the inference time taken to generate one complete response in seconds.
@@
Results. We present our main results in Figure 2 and additional ones in Tables 2, 3 and 4 in Appendix D. InferenceGuard achieves the highest safety rates with all models (reaching up to 94.46% on Alpaca, 98.45% on Llama-3.1-8B-Instruct, 98.97% on Vicuna, and 100% on Beaver), with minimal latency overhead in comparison to baselines (see also Table. 9). With Beaver, our method dominates the Patero front, achieving the highest rewards without any unsafe responses. Although Lagrangian methods can have a reasonable average cost, they fail to satisfy the safety constraints. Moreover, they are too safe on already safe answers, hindering their rewards. The RE-Control intervention method underperforms for all models in our setting. ARGS can provide safe answers but with very poor rewards because most answers are very short to avoid breaking the safety constraint. Among augmented safety methods, Best-of-N’s inability to leverage intermediate signals and beam search’s blindness to past mistakes lead to more unsafe answers and inferior performance compared to InferenceGuard.
@@
Figure 3 provides a better view of the reward and safety distributions. The figure shows that InferenceGuard consistently achieves higher rewards while maintaining low cumulative costs, outperforming other methods. Its cumulative cost stays just under the safety budget while maximizing the reward as suggested by our theoretical contributions. Finally, we observe that the trained critic helps to better guide the search on intermediate trajectories in both the unaligned Alpaca-7B model and the safety-aligned Beaver-7B model. We further provide ablation studies, limitations, latency overhead, critic training details, and qualitative comparisons of generated answers in Appendix D.
Figure 3: Reward and cost distributions of responses generated from Alpaca-7B (top) and Beaver-v3 (bottom) on PKU-SafeRLHF data. The left y-axis indicates the reward, while the right y-axis shows the cumulative cost. InferenceGuard outperforms baselines, both in terms of rewards and safety costs.